# WaveletGPT: Wavelet Inspired LLMs

### v2

@@ Abstract
Large Language Models (LLMs) have ushered in a new wave of artificial intelligence advancements impacting every scientific field and discipline. We live in a world where the data around us, e.g., text, audio, and music, has a multi-scale structure associated with it. This paper infuses LLMs with a traditional signal processing idea, wavelets, during pre-training to take advantage of the structure. Without adding any extra parameters to a GPT-style LLM architecture in academic setup, we achieve the same pre-training performance almost twice as fast in text, raw audio, and symbolic music. This is achieved by imposing a structure on intermediate embeddings. When trained for the same number of training steps, we achieve significant gains in performance, comparable to pre-training a larger neural architecture. Our architecture allows every next token prediction, access to intermediate embeddings at different temporal resolutions in every layer. This work will hopefully pave the way for incorporating multi-rate signal processing ideas into traditional LLM pre-training. Further, we showcase pushing model performance by improving internal structure instead of just going after scale. 111* This work was carried out in early 2023 while Prateek Verma was with the Stanford Natural Language Processing Group in the Department of Computer Science at Stanford University.
@@ 1 Introduction and Related Work
LLMs have ushered in a super-renaissance of AI advancements and are touching every scientific and engineering discipline. At the heart of this is the Transformer architecture (Vaswani et al., 2017), initially proposed for machine translation. Transformer architecture became the backbone of GPT (Generative Pretrained Transformer) language models (Brown et, 2020) first proposed by Open-AI. Modern LLMs are trained on a straightforward objective: To predict the next token given the previous context, preserving the causality. This not only works for language but also for robotics (Brohan et al., 2023b; a), protein sequences (Madani et al., 2020), raw audio waveforms(Verma & Chafe, 2021), acoustic and music tokens (Huang et al., 2019; Verma & Smith, 2020; Borsos et al., 2023), videos (Yan et al., 2021) etc. This simple recipe of tokenization/creating an embedding and feeding it to transformers also has given rise to non-causal architectures such as BERT(Devlin et al., 2019), Vision Transformers (Dosovitskiy et al., 2021), Audio Transformers (Verma & Berger, 2021) and Video Transformers (Selva et al., 2023). The recent surge in multi-modal LLMs similar to Gemini family (Team et al., 2023) or Chameleon Chameleon (2024) would pave the way computers able to reason like humans. With increased performance by scale, LLMs are reaching hundreds of billions of parameters (Brown et, 2020) with Google’s Switch Transformer even reaching trillion parameters (Fedus et al., 2022). Recent concerns suggest AI research is shifting from academia to industry, according to a Washington Post article by Nix (2024). The theme of this work is to enhance LLM capabilities to match those of larger architectures or achieve equivalent performance in fewer training steps. We extract intermediate embeddings from each decoder block and impose a hierarchical multi-scale structure without adding parameters. The signals across tokens in the intermediate layers are extracted, which we modify, similar to wavelet decomposition, while maintaining causality (Figure 1). Unlike previous techniques that enhance smaller models with larger ones, our approach focuses on improving performance during pre-training. A common approach is knowledge distillation Hinton et al. (2015), where a larger model guides a smaller one. Gu et al. (2024) used KL divergence to enhance next-token prediction from teacher model feedback, still relying on a powerful model rather than training the smaller one from scratch. A line of work, such as Nawrot et al. (2022), proposed hierarchical transformers using upsampling-downsampling operations, similar to the hourglass U-Net architecture Long et al. (2015) in computer vision. This approach achieves comparable results to Transformers but with more efficient computation. Clockwork RNN (Koutnik et al., 2014) improves long-context modeling by splitting RNN neurons into modules that update at different clock rates. Only a few modules activate at each time step, enabling efficient learning of long-term dependencies. In contrast, our approach modifies intermediate embeddings with simple tweaks, without using separate learning modules or varying update rates. Clockwork RNN (Koutnik et al., 2014) improves long-context modeling by splitting RNN neurons into modules that update at different clock rates. Only a few modules activate at each time step, enabling efficient learning of long-term dependencies. In contrast, our approach modifies intermediate embeddings with simple tweaks, without using separate learning modules or varying update rates. Model pruning (Sun et al., 2024) removes weights based on their salience, to match the same performance as a large model like LLAMA (Touvron et al., 2023), with fewer compute flops during inference. However, this approach still relies on starting with a pre-trained large model than training from scratch. We also exclude quantization methods like Dettmers et al. (2024), which focus on improving inference or fine-tuning existing models. The other line of work is tinkering with the intermediate embeddings. Tamkin et. al (2020) proposed hand-tuned filters on the Discrete Cosine Transform- DCT over the entire context length (Ahmed et al., 1974) of the latent space for different NLP tasks for non-causal BERT (Devlin et al., 2019), making them not applicable for causal applications such as language modeling. There have been work on applying ideas from signal processing-like methods to BERT-like non-causal architectures. We discuss two here, FNet and WavSPA. They focus on improving attention, which is different from our work on GPT which retains vanilla attention layer. FNet proposed by Lee-Thorp et al. (2022) removes the costly attention mechanism replacing with a 2-D FFT block. This operation is non-causal as it looks into future tokens for computing 2-D FFT. WavSpA (Zhuang et al., 2024) carries attention mechanism in the wavelet space. Since the wavelet transform is a multi-resolution, captureing long-term dependencies at various time scales, the input sequences are transformed into wavelet space, and the attention mechanism is carried out and then reconstructed. However, computing wavelet transform is non-causal, making them non applicable for GPT based LLMs as they look at the entire sequence length for capturing variations from coarsest to finest scales (as can be seen in Figure 1 of (Zhuang et al., 2024)). This paper modifies only the intermediate embeddings of a LLM model. Our work is inspired by neuroscience, which provides evidence that the human brain learns multi-scale representations for language at multiple time scales (Caucheteux et al., 2023) instead of fixed resolution representations. Our paper explicitly imposes multi-scale representation onto every intermediate decoder embedding at different dimensions. The contribution of the paper is as follows:
@@
  • We propose, to the best of our knowledge, the first instance of incorporating wavelets into LLM pretraining. We add multi-scale filters onto each of the intermediate embeddings of decoder layers using Haar/learnable wavelet pipeline. This allows every next token prediction access to multi-scale intermediate embeddings in every decoder layer instead of fixed-resolution representations.
  • We show to speed the pre-training of a shrunk down GPT like transformer-based LLM in the range of 40-60%, with adding multi-scale structure. With the same number of training steps, the model gives a substantial non-trivial performance boost, akin to adding several layers or parameters.
Figure 1: Manipulating signals in between decoder blocks of GPT. For each of these signals we compute a 1-D causal discrete haar wavelet transform or learnable approximation at different levels to mimic the multi-scale structure that exists for text, raw audio, symbolic music. Fig on right is from Gao & Yan (2006), which gives a detailed account of non-stationary signal processing for 1-D signals. We take the leftmost route of approximate coefficients, which allows us coarsest to finest scales.
@@ 2 Dataset
We use three open-source datasets from three different domains: natural language, symbolic music, and raw audio waveform. For text, we choose text-8 (Mikolov et al., 2012). We choose this over other datasets as i)it is popular and widely cited character-level language modeling dataset and ii) use a simple vocabulary (space + 26 lowercase characters) to detach the effects of various tokenizers. It has 100M characters with split training split as given by Al-Rfou et al. (2019). For raw audio, the goal is predicting the next sample given context. We use the YouTube-Mix-8 dataset, used for long-context modeling (Goel et al., 2022; Verma, 2022). Our vocabulary size is 256, with a sampling rate 16KHz as input is 8-bit. We use a third dataset, MAESTRO (Hawthorne et al., 2019), containing over 1000 MIDI files of classical music pieces. We use tokenizer proposed by Huang et al. (2019), which converts MIDI tracks into discrete tokens with a vocabulary size 388. The goal in all three modalities is not to chase state-of-the-art performance, as this paper was written in an academic setting with very few computational resources. We compare pre-training performance to the shrunk-down version of GPT with/without adding multi-scale structure to the embeddings using Haar or learnable kernels.
@@ 3 Methodology
This section will describe the approach to incorporating wavelets into transformer-based Large Language models while retaining the causality assumption. The ideas described here are generic and can be easily extrapolated to setups without a Transformer architecture e.g. state space architectures.
3.1 Incorporating Wavelets into Intermediate Embeddings
For any signal, we compute a version of the discrete wavelet transform and incorporate it back into the signal. Let $x_{(i)}^{l}$ be the output of the $l^{th}$ decoder layer, representing the activation along the $i^{th}$ coordinate, with a dimension equal to the context length $L$ of the transformer-based GPT model. In the original GPT architecture with $N+1$ layers and embedding dimension $E$, we obtain $N\cdot E$ signals of length $L$ from intermediate embeddings between decoder blocks, where $E$ ranges from $[0-128)$ dimensions. A wavelet is a signal with zero mean and non-zero norm, designed to address the limitations of traditional Fourier-based representation. For any signal $x[n]$, the discrete wavelet transform resembles passing the signal through filters of varying resolutions, as illustrated in Figure 2. We will use the Haar wavelet, a family of square-shaped functions, throughout this paper, obtained from a mother wavelet via scaling and shifting operations. Given a mother wavelet function $\psi$, we come up with the child wavelets as $\psi_{j,k}[n]$, where $j$ is the scaling factor and $k$ the shift factor.
@@ 3.1 Incorporating Wavelets into Intermediate Embeddings
These signals are shifted and scaled to capture information at various time scales, with $n$ representing time or the context length. This concept resembles the diagram in Figure 1, which illustrates capturing different signals in the intermediate layers of Transformer decoders at various resolutions. We now define the discrete wavelet transform, which passes any signal through filters and downsampling operations. This process, shown in Figure 2, is similar to a convolutional neural network (CNN) like ResNet (He et. al, 2016), featuring learned convolutional filters analogous to $h[n]$ and $g[n]$, along with downsampling, such as max pooling. In traditional convolutional architectures, we typically follow one branch of Figure 2, recursively taking the output of filters and downsampling. This similarity contributed to the popularity of wavelets in the 1990s and 2000s for image understanding, reflecting parallels with convolutional architectures (Huang & Aviyente, 2008; Kingsbury & Magarey, 1998). As we use Haar wavelets, this involves passing the signal through low-pass and high-pass filters corresponding to the kernels $g[n]$ and $h[n]$. The Haar wavelet transform averages and computes differences, with impulse responses $g[n]=\left[\frac{1}{2},\frac{1}{2}\right]$ and $h[n]=\left[\frac{1}{2},-\frac{1}{2}\right]$. Figure 2 provides a detailed explanation of the discrete wavelet transform. For a 1-D signal $x[n]$ of length $L$, we obtain level 1 coefficients by applying filters $g[n]$ and $h[n]$, followed by downsampling. Thus, the approximation coefficients $y_{approx}$ and $y_{detail}$ result from an LTI system defined by convolution followed by downsampling by two, seen in Equation 2. This behavior is reflected in convolution in Algorithm 1.
$y_{\text{approx}}[n]=\sum_{k=-\infty}^{\infty}x[k]g[2n-k]$ ; $y_{\text{detail}}[n]=\sum_{k=-\infty}^{\infty}x[k]h[2n-k]$ (2)
@@
To obtain multi-scale representations of the original signal, the operation for $x[n]$ is recursively applied to $y_{a}$ (approx) to derive level 2 wavelet coefficients $y^{2}_{a}$ and $y^{2}_{d}$ (detail). Here, $x[n]$ represents intermediate signals across the context length at each decoder block output in the LLM. The approximate coefficients $y_{a}$ and $y_{d}$, along with their decompositions $\{y_{a},y_{d},y^{2}_{a},y^{3}_{a},y^{4}_{a},\ldots\}$, are used for further processing. Notably, $y^{2}_{a},y^{3}_{a},y^{4}_{a}$ have lengths reduced by factors of $2,4,8,\ldots$. The Haar wavelet transform averages adjacent samples while preserving causality by averaging current and past samples. Higher-order coefficients capture averages over larger context lengths, as shown in Figure 2. We can continue until only a single scalar value remains, representing the mean of the signal. The Haar wavelet transform computes averages and differences to create a multi-resolution representation, capturing low and high frequencies at different resolutions. Figure 2 illustrates the same signal captured at coarser and finer representations using Haar wavelets, applied to intermediate embeddings, allowing each next token prediction access to these representations. For the case of learnable wavelet kernels, we create a multi-resolution representation by varying the kernel size (Algorithm 1) to allow the LLM to learn the optimal kernels optimized for next token prediction.
Figure 2: (Bottom L): A tree structure depicting 3-level filter bank that gives us a signal at different resolutions. We get approximate coefficients by passing it through an impulse response corresponding to the chosen wavelet and recursively down-sampling it. (Top left) Computing the approximate and detailed coefficients at various levels. We recursively take first-order averages/differences followed by downsampling until we get only a single scalar representative of the input signal. (Right) For a 32-length signal different levels of approximate coefficients of haar wavelet capture the signal from the coarsest to finest. The figure on left are redrawn from (Flores-Mangas, 2014). (R) We compute embeddings moving at different rates via causal wavelet approximation, where certain embedding dimensions evolve at the coarsest level (similar to level 5), while others follow a finer resolution (level 2). This infuses multi-scale information in all embeddings for decoder layers for every token.
@@ 3.2 Connecting wavelets and LLM embeddings
In many signal processing applications, first-order detail coefficients and approximate coefficients help understand signals at various levels. We aim to do the same but with signals from intermediate transformer embeddings across tokens. However, we focus only on approximate coefficients. Our premise is that real-world data is structured—text ranges from letters to words, sentences, and topics, while symbolic music ranges from notes to motifs and pieces. Using the Haar wavelet, this can be approximated as a simple averaging operation, as described earlier. For the learnable version, we allow weights of the kernel for multi-scale version to be optimized according to how best we can predict the next token. Continuing with the approximate coefficients will eventually yield a single scalar, the average of the entire signal in the case of the Haar wavelet. To match the original signal’s sequence length from the approximation coefficients, several methods can be employed, including up-sampling. For clarity, we refer to the signal approximated at a specific level with the same length as the ”approximate signal” at that level, distinguishing it from the shorter approximate coefficients. In Figure 2 (R), to obtain the signal approximation at various levels matching the original input signal $x[n]$, we apply the wavelet kernel by multiplying the approximate coefficients with the kernel for that level (e.g., $[1,1]$, $[1,1,1,1]$, etc.). This is illustrated in the piecewise constant function shown in Figure 2. Different LLM embedding coordinates define unique resolution kernels, each corresponding to a specific scale for data capture. The reconstructed signal $x_{\text{recon}}[n]$, a method to derive the *approximate signal*, is computed from wavelet coefficients $c_{j}$ at level $j$ as:
$\vspace{-0.15cm}x^{j}_{\text{recon}}[n]=\sum_{k}c_{k}\cdot\psi_{j,k}[n]\vspace{-0.15cm}$ (3)
@@
Equation 3 requires storing child wavelets at various approximations, complicating the process and rendering it non-causal as computing $c_{k}$ takes into account the entire signal. Due to the dependence of $c_{k}$ to future information, we cannot use this to reconstruct the signal from its approximate coefficients. To adapt this for LLM we simplify the computation of the *approximate signal* in a differentiable manner using a variant of the equation from Equation 3 in both multi-resolution learnable/non-learnable kernels settings. For the Haar wavelet, we compute a average of the input signal with varying kernel lengths, increasing the length until it approximates the entire signal. The kernel length determines the level of signal approximation. LLMs operate under a causality assumption, modifying the signal at a location using prior samples within the kernel length. We zero-pad the signal to the left when window length is shorter than kernel. Wavelet transform at different levels gives several versions of the signal at different resolution which can mess up the structure of the intermediate Transformer embeddings. To address this, we create different resolutions for signal approximations parameterized by the embedding dimension. In Section 4.4, we make these kernels learnable, allowing the architecture to maintain multi-scale operation (Equation 3), with learnable weights with $x_{\text{recon}}[n]$ now being learned. The resolution is parameterized by the embedding coordinate is described next.
Algorithm 1 Wavelet-GPT
$E$: Model or Embedding Dimension
$L$: Context Length
$N+1$: Number of Decoder Layers
for layer $l=1,2,\ldots,N$ do
     $\mathbf{x}^{l}\leftarrow\text{Output of Transformer $l^{th}$ Decoder Block}$ //Dimension $E$ x $L$
     $\mathbf{xn}^{l}\leftarrow\text{Modified Transformer Embedding Replacing $\mathbf{x}^{l}$}$
     $\mathbf{xn}^{l}_{(i)}\leftarrow\mathbf{x}^{l}_{(i)}\quad\text{For Embedding dimension \quad}i>E/2$
     $\mathbf{f(i)}\leftarrow 2^{F}\text{where}$ //Finding kernel length function of embedding coordinate nearest power of 2
     $\quad\quad\quad\quad F=int(L_{k}*(i-E/2)/(E/2-1))\quad L_{k}=\lfloor\log_{2}(L)\rfloor+1\quad\text{i}<=E/2$
     $\mathbf{xn_{(i)}^{l}(k)}\leftarrow\mathbf{\frac{1}{f(i)}\sum_{m=k-f(i)}^{k}x_{(i)}^{l}(m)}\quad\text{i}<=E/2$ // For Non-learnable fixed Haar wavelet
     $\mathbf{xn_{(i)}^{l}(k)}\leftarrow\mathbf{\sum_{m=0}^{f(i)-1}h(m)\cdot x_{(i)}^{l}(k-m)}\quad\text{i}<=E/2$ // For learnable wavelet kernel $h$
end for
3.3 Wavelet Coefficients by Embedding Dimension Coordinates
One option is to compute the approximate signals for each coordinate signal $x_{(i)}^{l}$ across all decoder layers at levels I to IX. For a context length of 512, this would require nine additional signals with resolutions of 512, 256, 128, 64, 32, 16, 8, 4, and 2, significantly increasing the architecture’s complexity and necessitating major modifications to our GPT model. To address this, we propose a novel solution: instead of computing all levels of approximate signals for every intermediate embedding dimension, we parameterize the level by the embedding dimension index. We want to steer the embeddings only a little into the inductive biases we impose to avoid too much tinkering with that they learn. Transformers have been wildly successful without incorporating any inductive biases. Ideally, we want the best of both worlds, nudging intermediate GPT embeddings in only half of the dimensions. We adjust intermediate GPT embeddings in only half the dimensions. Embeddings from $E/2$ to $E$ (coordinates 64 to 128 when $E=128$) remain unchanged. For the rest, we apply processing based on their index $i$. Mathematically, if $x^{l}{(i)}$ is an intermediate embedding after the $l^{th}$ decoder layer along the $i^{th}$ dimension, the modified signal $xn^{l}{(i)}$ equals $x^{l}_{(i)}$ for $i\in[E/2,E]$. For $20\leq i<E/$ 2, we impose structure using an approximate signal, calculated from wavelet coefficients corresponding to the index $i$. We use a mapping function $f$ that takes coordinate $i$ (ranging from 00 to $E/2$) and returns the kernel size corresponding to approximation levels from $I$ to $IX$. The linear function gradually increases from level $I$ (kernel size 2 at $i=0$) to level $IX$ (kernel size 512 at $i=E/2$, or the coarsest representation for a generic case, i.e., a scalar). Now, let us find out how we compute the modified new signal $xn^{l}_{(i)}$ that replaces the original intermediate Transformer embeddings $x^{l}_{(i)}$. $f(i)$ denotes the kernel size for the coordinate $i$. Now, the modified signal is:
$xn_{(i)}^{l}=x_{(i)}^{l}\text{ for }i>E/2$ ; $xn_{(i)}^{l}(k)=\frac{1}{f(i)}\sum_{m=k-f(i)}^{k}x_{(i)}^{l}(m).$ (4)
For cases where $k-f(i)<0$, we zero-pad the signal to ensure valid average/kernel computation. Specifically, for the Haar wavelet, the modified signal acts as a causal moving average filter with finite length, averaging the embedding signal along the $i^{\text{th}}$ coordinate with a kernel size determined by $f(i)$. This operation does not introduce new parameters, maintaining causality in LLMs and preventing future token leakage as seen in Equation 4. We can extend this approach to learn an optimal kernel specific to the task. In Algorithm 1, each value of the modified signal at token $k$ is computed using a convolution with a learned kernel $h(.)$ and variable length $f(i)$, parameterized by the embdding coordinate dimension $i$. Each kernel is learned independently for every signal in LLM.

### v4

# Wavelet GPT - Wavelet Inspired Large Language Models

@@ Abstract
Large Language Models (LLMs) have ushered in a new wave of artificial intelligence advancements impacting every scientific field and discipline. We live in a world where most of the data around us, e.g., text, audio, and music, has a multi-scale structure. This paper infuses LLMs with a traditional signal processing idea, namely wavelets, during pre-training to take advantage of the structure. Without adding any extra parameters to a GPT-style LLM architecture in an academic setup, we achieve the same pre-training performance almost twice as fast in text, audio, and images. This is done by imposing a structure on intermediate embeddings. When trained for the same number of training steps, we achieve significant gains in performance, which is comparable to pre-training a larger neural architecture. Further, we show this extends to the Long Range Arena benchmark and several input representations such as characters, BPE tokens, bytes, waveform, math expression, and image pixels. Our architecture allows every next token prediction access to intermediate embeddings at different temporal resolutions in every decoder block. We hope this will pave the way for incorporating multi-rate signal processing into pre-training instead of going after scale.
Machine Learning, ICML
@@ 1 Introduction and Related Work
LLMs have ushered in a super-renaissance of AI advancements and are touching every scientific and engineering discipline. At the heart of this is the Transformer architecture (Vaswani et al., 2017), initially proposed for machine translation. Transformer architecture became the backbone of GPT (Generative Pretrained Transformer) language models (Brown et, 2020) first proposed by Open-AI. Modern LLMs are trained on a straightforward objective: To predict the next token given the previous context, preserving the causality. This not only works for language but also for robotics (Brohan et al., 2023b, a), protein sequences (Madani et al., 2020), raw audio waveforms(Verma & Chafe, 2021), acoustic/music tokens (Huang et al., 2019; Verma & Smith, 2020; Borsos et al., 2023), videos (Yan et al., 2021) etc. This simple recipe of tokenization/creating an embedding and feeding it to transformers also has given rise to non-causal architectures such as BERT(Devlin et al., 2019), Vision Transformers (Dosovitskiy et al., 2021), Audio Transformers (Verma & Berger, 2021) and Video Transformers (Selva et al., 2023). With increased performance by scale, LLMs are reaching hundreds of billions to trillions of parameters (Brown et, 2020; Fedus et al., 2022).
Recent concerns suggest AI research is shifting from academia to industry, according to a Washington Post article by (Nix, 2024). This work aims to enhance LLM capabilities to match those of larger architectures or achieve equivalent performance in fewer training steps. Knowledge distillation (Hinton et al., 2015), uses a larger model to guide a smaller one. (Gu et al., 2024) used KL divergence to enhance next-token prediction from teacher model feedback model rather than training the smaller one from scratch. Model pruning (Sun et al., 2024) removes weights to match the same performance as a large model like LLAMA (Touvron et al., 2023), with fewer compute flops during inference, still relying on a larger model. (Dettmers et al., 2024) focus on improving inference or fine-tuning existing models. Unlike distillation and pruning our approach focuses on improving performance during pre-training from scratch. (Nawrot et al., 2022), proposed hierarchical transformers using upsampling-downsampling operations achieve results comparable to those of Transformers but with more efficient computation. Clockwork RNN (Koutnik et al., 2014) improves long-context modelling by splitting RNN neurons into modules that update at different clock rates. Only a few modules activate at each time step. Our approach modifies intermediate embeddings with simple tweaks without using separate learning modules or varying update rates.
Tinkering with the intermediate embeddings: (Tamkin et. al, 2020) proposed hand-tuned filters on the Discrete Cosine Transform- DCT (Ahmed et al., 1974) of the latent space for different NLP tasks for non-causal BERT (Devlin et al., 2019). Computing DCT over context length makes it not applicable for causal applications such as language modelling. There has been work on applying ideas from signal processing-like methods to BERT-like non-causal architectures. We discuss two here, FNet and WavSPA. They focus on improving attention block, which differs from our work on GPT, which retains a vanilla attention layer. FNet proposed by (Lee-Thorp et al., 2022) removes the costly attention mechanism, replacing it with a 2-D FFT block. This operation is non-causal as it looks into future tokens for computing 2-D FFT. WavSpA (Zhuang et al., 2024) carries attention mechanism in the wavelet space. The input sequences are transformed into wavelet space, and the attention mechanism is carried out and then reconstructed. However, computing wavelet transform is non-causal, making them non-applicable for GPT-based LLMs as they look at the entire sequence length (Fig 1 (Zhuang et al., 2024)).
@@
Our work is inspired by neuroscience, which provides evidence that human brain learns multi-scale representations for language at multiple time scales (Caucheteux et al., 2023) instead of fixed-resolution representations. We impose multi-scale representation onto every intermediate decoder embedding at different dimensions. To the best of our knowledge, the paper’s contributions are: 1) We propose the first instance of incorporating wavelets into LLM pre-training. We add multi-scale filters onto each of the intermediate embeddings of decoder layers using the Haar/learnable wavelet pipeline. This allows every next token prediction access to multi-scale intermediate embeddings instead of being fixed-resolution in every decoder layer representation. 2) We show speedups in pre-training of GPT, like transformer-based LLM in the range of 40-60%, with adding a multi-scale structure. With same training steps, the model gives a performance boost akin to adding several layers.
@@ 2 Dataset
We use four open-source datasets from four domains for next token prediction: natural language, symbolic music, speech tokens, and raw audio waveform. For text, we choose text-8 (Mikolov et al., 2012). We choose this over other datasets as i)it is a famous and widely cited character-level language modelling dataset, and ii) it uses a simple vocabulary (space + 26 lowercase characters) to detach the effects of various tokenizers. It has 100M characters with split training split as given by (Al-Rfou et al., 2019). For raw audio, the goal is to predict the next sample given the context. We use the YouTube-Mix-8 dataset for long-context modeling (Goel et al., 2022; Verma, 2022). Our vocabulary size is 256, with a sampling rate 16KHz as input is 8-bit. We use a third dataset, MAESTRO (Hawthorne et al., 2019), containing over 1000 MIDI files of classical music pieces with a tokenizer proposed by (Huang et al., 2019), which converts MIDI tracks into discrete tokens with a vocabulary size of 388. Finally, we use 1000 hours of LibriSpeech dataset and a widely used ENCODEC (Défossez et al., 2022) tokenizer in a setup similar to VALL-E (Wang et al., 2023) to model acoustic tokens  [^1] The goal in this work, is how well we can model the coarsest tokens, as errors in modelling the coarser tokens will lead to the finer tokens being modelled incorrectly as they are conditioned on the coarsest token, thus affecting model performance. The goal in all four modalities is not to chase state-of-the-art pre-training performance, as this paper was written in an academic setting with very few computational resources. We show how the model performs in pre-training instead of post-training, as the goal is to build better foundational architectures with the same parameters, pushing the capabilities of smaller decoder architectures.
Figure 1: Manipulating signals between GPT decoder blocks by computing 1-D causal discrete haar wavelet transform/learnable approximation at different levels capturing multi-scale structure for each signal. (Right) From (Gao & Yan, 2006) explaining non-stationary signal processing for signals. Leftmost route of approximate coefficients to model coarsest to finest scales.
Figure 2: (Bottom L): A 3-level filter bank tree generates signals at different resolutions. Approximate coefficients are computed by applying a wavelet’s impulse response & recursively down-sampling. (Top L): Approximate and detailed coefficients are iteratively calculated via first-order averages/differences and down-sampling until a single scalar represents the signal. (R): For a 32-length signal, Haar wavelet captures coarsest to finest approximations and is redrawn from (Flores-Mangas, 2014). Embeddings evolve at different rates via causal wavelet approximation, with coarse (level 5) and fine (level 2) resolutions, embedding multi-scale information into decoder layers for every token.
@@ 3 Methodology
This section will describe the approach to incorporating wavelet inspired computation into transformer-based Large Language Models while retaining causality. The ideas described here are generic as we tinker intermediate embeddings. Thus theycan be easily extrapolated to non-Transformer architectures, e.g. state space model. For any GPT signal, we compute a version of the discrete wavelet transform and incorporate it back into the signal. Let $x_{(i)}^{l}$ be the output of the $l^{th}$ decoder layer, representing the activation along the $i^{th}$ coordinate, with a dimension equal to the context length $L$ of the transformer-based GPT model. In the original GPT architecture with $N+1$ layers and embedding dimension $E$, we obtain $N\cdot E$ signals of length $L$ from intermediate embeddings between decoder blocks, where $E$ ranges from $[0-128)$ dimensions. For any signal $x[n]$, the discrete wavelet transform resembles passing the signal through filters of varying resolutions, as illustrated in Figure 2. We will use the Haar wavelet, a family of square-shaped functions this paper obtained from a mother wavelet via scaling and shifting operations. Given a mother wavelet function $\psi$, the child wavelets as $\psi_{j,k}[n]$, where $j$ is the scaling factor and $k$ is the shift factor.
@@
These signals are shifted and scaled to capture information at various time scales, with $n$ representing time or the context length. This concept resembles the diagram in Figure 1, which illustrates capturing different signals in the intermediate layers of Transformer decoders at various resolutions. Discrete wavelet transform, which passes any signal through filters and downsampling operations. This process, shown in Figure 2, is similar to a convolutional neural network (CNN) like ResNet (He et. al, 2016), featuring learned convolutional filters analogous to $h[n]$ and $g[n]$, along with downsampling, such as max pooling. In convolutional architectures, we follow one branch of Figure 2, recursively taking the output of filters and downsampling. This similarity contributed to popularity of wavelets in the 1990/2000s for image understanding, reflecting parallels with convolutional architectures (Huang & Aviyente, 2008; Kingsbury & Magarey, 1998). For Haar wavelets, this is passing the signal through low-pass and high-pass filters corresponding to the kernels $g[n]$ and $h[n]$. The Haar wavelet transform averages and computes differences, with impulse responses $g[n]=\left[\frac{1}{2},\frac{1}{2}\right]$ and $h[n]=\left[\frac{1}{2},-\frac{1}{2}\right]$. Figure 2 provides a detailed explanation of the discrete wavelet transform. For a 1-D signal $x[n]$ of length $L$, we get level 1 coefficients by filters $g[n]$ and $h[n]$, followed by downsampling. Thus, the approximation coefficients $y_{\text{approx}}$ and $y_{\text{detail}}$ result from a LTI system defined by convolution followed by downsampling by two (Equation 2). This is seen in Algorithm 1 with $\text{type}\in\{\text{approx},\text{detail}\}$ and $f_{\text{approx}}=g$, $f_{\text{detail}}=h$.
$y_{\text{type}}[n]=\sum_{k=-\infty}^{\infty}x[k]\,f_{\text{type}}[2n-k]$ (2)
@@
To obtain multi-scale representations of the original signal, the operation for $x[n]$ is recursively applied to $y_{a}$ (approx) to derive level 2 wavelet coefficients $y^{2}_{a}$ and $y^{2}_{d}$ (detail). Here, $x[n]$ represents intermediate signals across the context length at each decoder block output in the LLM. The approximate coefficients $y_{a}$ and $y_{d}$, along with their decompositions $\{y_{a},y_{d},y^{2}_{a},y^{3}_{a},y^{4}_{a},\ldots\}$, are used for further processing. Notably, $y^{2}_{a},y^{3}_{a},y^{4}_{a}$ have lengths reduced by factors of $2,4,8,\ldots$. The Haar wavelet transform averages adjacent samples while preserving causality by averaging current and past samples. Higher-order coefficients capture averages over larger context lengths, as shown in Figure 2. We can continue until only a single scalar value remains, representing the mean of the signal. The Haar wavelet transform computes averages and differences to create a multi-resolution representation, capturing low and high frequencies at different resolutions. Figure 2 illustrates the same signal captured at coarser and finer representations using Haar wavelets, applied to intermediate embeddings, allowing each next token prediction access to these representations. For the case of learnable wavelet kernels, we create a multi-resolution representation by varying the kernel size (Algorithm 1) to allow the LLM to learn the optimal kernels optimized for the next token prediction.
@@ 3.1 Connecting wavelets and LLM embeddings
In many signal processing applications, first order detailed along with approximate coefficients captures signals at various levels. We do the same with signals from intermediate transformer embeddings across tokens. Our premise is that real-world data is structured: text ranges from letters to words, sentences, and topics, symbolic music ranges from notes, motifs and pieces with speech from phonemes, morphemes to syllables to phrases. With Haar wavelet, this can be approximated as a simple averaging operation. For the learnable version, the weights of the kernel for the multi-scale version are optimized to predict the next token. Continuing with the approximate coefficients will eventually yield a single scalar, the average of the entire signal in the case of the Haar wavelet. Several methods can be employed to match the original signal’s sequence length from the approximation coefficients, including up-sampling. We refer to the signal approximated at a specific level with the same length as the ”approximate signal” distinguishing it from shorter approximate coefficients. In Figure 2 (R), to obtain the signal approximation at various levels matching the original input signal $x[n]$, we apply the wavelet kernel by multiplying the approximate coefficients with the kernel for that level (e.g., $[1,1]$, $[1,1,1,1]$, etc.). This piecewise constant function is shown in Figure 2. LLM embedding coordinates define unique resolution kernels, each corresponding to a specific scale of data. The reconstructed signal $x_{\text{recon}}[n]$, a method to derive the *approximate signal*, is computed from wavelet coefficients $c_{j}$ at level $j$ as:
$x^{j}_{\text{recon}}[n]=\sum_{k}c_{k}\cdot\psi_{j,k}[n]$ (3)
@@
Equation 3 requires storing child wavelets at various approximations, complicating the process and rendering it non-causal as computing $c_{k}$ considers the entire signal. Due to the dependence of $c_{k}$ on future information, we cannot use this to reconstruct the signal from its approximate coefficients. To adapt this for LLM, we simplify the computation of the *approximate signal* in a differentiable manner using a variant from Equation 3 in both multi-resolution learnable/non-learnable kernel settings. For the Haar wavelet, we compute an average of the input signal with varying kernel lengths, increasing the length until it approximates the entire signal. The kernel length determines the level of signal approximation. LLMs operate under a causality assumption, modifying the signal at a location using prior samples within the kernel length. We zero-pad the signal to the left when the window length is shorter than the kernel. Wavelet transform at different levels gives several versions of the signal at different resolutions, which can mess up the structure of the intermediate embeddings. To address this, we create different resolutions for signal approximations parameterized by embedding dimension. In Section 4.4, we make these kernels learnable, allowing the architecture to maintain multi-scale operation (Equation 3), with learnable weights with $x_{\text{recon}}[n]$ now being learned with resolution parameterized by latent dimension.